You measure 27 textbooks' weights, and find they have a mean weight of 31 ounces. Assume the population standard deviation is 8.7 ounces. Based on this, construct a 95% confidence interval for the true population mean textbook weight. Give your answers as decimals, to two places < μ μ < License
Answer #1

Solution :

Given that,

Sample size = n = 27

Z/2 = 1.96

Margin of error = E = Z/2* ( /n)

= 1.96 * (8.7 / 27)

Margin of error = E = 3.28

At 95% confidence interval estimate of the population mean is,

- E < < + E

31 - 3.28 < < 31 + 3.28

27.72 < < 34.28